OpenStudy (anonymous):

Does Y-X/2 = 1/2*(Y-X) ? Does 1/2*(Y-X) = Y/2 - X/2 ??

OpenStudy (anonymous):

Yes for the first one if Y-X/2 is actually with parenthesis (Y-X)/2 Yes for second, because for example lets have y=2 and x=2 (2-2)/2 = 0 is the same as 2/2 - 2/2 = 0

OpenStudy (anonymous):